OpenStudy (anonymous):

Complete the ordered pair (1/8, ) so that it is a solution to the equation: 32x + 2y = 26

OpenStudy (johnweldon1993):

well we have an x value right? (x,y) so we can plug that in for x in the equation 32 ( 1/8) + 2y = 26 what is 32 times 1/8? you do 32 times 1......divided by 8....also known as 4 so we have 4 + 2y = 26 subtract 4 from both sides 2y = 22 divide both sides by 2 y = 11 test it out. 32(1/8) + 2(11) = 26 4 + 22 = 26? 26 = 26? yes it does...so that is your answer :)

OpenStudy (anonymous):

oh..i see... thank you.

OpenStudy (anonymous):

Complete the ordered pair ( , 18) so that it is a solution to the equation: 3x + 2y = 72 x is 12?

OpenStudy (johnweldon1993):

not a problem :) and..you are correct :)

OpenStudy (anonymous):

ok thank you.

OpenStudy (johnweldon1993):

anytime :)